### v2

@@ 6 Conclusion
Under the framework implied by the proposed definition of nominal outlyingness, we introduce a method for quantifying outlyingness for nominal data. Our proposal includes ideas from the association rule mining literature and can be seen as a extension of the work of [24]. We present a strategy that can be used to determine values for the hyperparameters that are involved in the calculation of the score of nominal outlyingness and provide details on support-based pruning, which seeks to reduce computational complexity. Furthermore, we define the concept of nominal outlyingness depth and the contribution matrix. These can be used to assess the source of nominal outlyingness by providing the user with the average length of highly frequent/infrequent sequences included in an observation, as well as with the extent to which each nominal variable has contributed to the score of outlyingness of an observation.
@@
Our proposed framework has been compared to two state-of-the-art algorithms for nominal outlyingness quantification, namely FPOF and FPI. The efficacy our method was illustrated via a series of simulations on 2700 synthetic data sets, where the results aligned with what was expected given the data generating process. In contrast, the two competing algorithms acted in a completely agnostic manner, inflating the scores for observations with nominal levels that were not genuinely outlying but were set by default to appear fewer times than others. We have further applied our method on four real-world data sets. Two of these data sets have not been previously used for testing outlier detection but rather for classification purposes, while the other two typically serve as benchmarking data sets for newly developed outlier detection algorithms. Our method has produced comparable and in certain instances even better results to these of FPOF and FPI, under the assumption of the minority class being the class of outliers in these data sets. Moreover, our proposed framework has selected suitable hyperparameter values without requiring user input, a feature that is not supported by the other two methods.
@@
The proposed framework has yielded some remarkable results that can be easily interpreted (thanks to the introduction of the nominal outlyingness depth and the contribution matrix) on the real-world data. However, it does not come with no limitations. A shortcoming of the framework presented is the way the hyperparameter $\operatorname*{\operatorname{MAXLEN}}$ is computed. For instance, if we are interested in highly infrequent itemsets and we have a large number of nominal variables with highly imbalanced level proportions, the expected number of sequences containing only the most frequent levels is typically large enough to produce sparse expected contingency tables. A direct consequence of this is that $\operatorname*{\operatorname{MAXLEN}}$ will be very large, leading to potentially redundant calculations and therefore an increased computational cost. This has already been observed in some cases in our simulations, where the depth was much lower than the computed value of $\operatorname*{\operatorname{MAXLEN}}$, indicating that sequences of greater length have probably had limited to no contribution to the obtained scores of nominal outlyingness. An amendment in the way $\operatorname*{\operatorname{MAXLEN}}$ is determined would alleviate this issue and reduce the computational complexity, without changing the way the results are interpreted and the conclusions derived. Additionally, one could extend the simulation study by enforcing a dependence structure among variables which are likely to be associated in a specific manner. This way, we can get rid of the independence assumption (which is typically not valid) and potentially obtain more interesting results.
@@
We conclude by giving some recommendations on how our framework could be used for extending the research that is currently being conducted in certain fields. The proposed framework could be used for assigning weights to observations in a data set prior to running a classification algorithm. As an example, one would perhaps want to downweigh subjects with a large score of nominal outlyingness, as these can be seen as potential outliers which may distort the classification output. This can be easily done by modifying the relevant loss function to account for observation weights. Similarly, one might wish to implement such an approach for clustering categorical data or deriving a version of the Trimmed K-Means [16] algorithm for nominal data. Finally, the proposed framework could contribute in the development of more reliable predictions in linear regression with nominal predictors. Extremities are commonly the reason behind biased parameter estimates and misleading predictions of a regression model; being able to diminish the impact of potential outliers could help mitigate these problems and develop robust regression algorithms for nominal data. These recommendations could pave the way for future research opportunities and culminate in the development of novel methodologies which will in turn provide insight on research questions of interest.
@@
The R code for the proposed method and the simulations conducted is available on GitHub by accessing the repository in: https://anonymous.4open.science/r/SONO-7CD4/.

### v3

@@ 6 Conclusion
Under the framework implied by the proposed definition of nominal outlyingness, we introduce a method for quantifying outlyingness for nominal data. Our proposal includes ideas from the association rule mining literature and can be seen as an extension of the work of [25]. We present a strategy that can be used to determine values for the hyperparameters that are involved in the calculation of the score of nominal outlyingness and provide details on support-based pruning, which seeks to reduce computational complexity. Furthermore, we define the concept of nominal outlyingness depth and the contribution matrix. These can be used to assess the source of nominal outlyingness by providing the user with the average length of highly frequent/infrequent sequences included in an observation, as well as with the extent to which each nominal variable has contributed to the score of outlyingness of an observation.
@@
Our proposed framework is compared to two state-of-the-art algorithms for nominal outlyingness quantification, namely FPOF and FPI. The efficacy our method is illustrated via a series of simulations on 2700 synthetic data sets, where the results align with what is expected given the data generating process. In contrast, the two competing algorithms act in a completely agnostic manner, inflating the scores for observations with nominal levels that are not genuinely outlying but are set by default to appear fewer times than others. We apply our method on four publicly available data sets. Two of these data sets have not been previously used for testing outlier detection but rather for classification purposes, while the other two typically serve as benchmarking data sets for newly developed outlier detection algorithms. Our method produces comparable and in certain instances even better results to these of FPOF and FPI, under the assumption of the minority class being the class of outliers in these data sets. Moreover, our proposed framework selects suitable hyperparameter values without requiring user input, a feature that is not supported by the other two methods.
@@
The proposed framework yields some remarkable results that can be easily interpreted (thanks to the introduction of the nominal outlyingness depth and the contribution matrix) on the publicly available data. However, it does not come with no limitations. A shortcoming of the framework presented is the way the hyperparameter $\operatorname*{L_{\text{max}}}$ is computed. For instance, if we are interested in highly infrequent itemsets and we have a large number of nominal variables with highly imbalanced level proportions, the expected number of sequences containing only the most frequent levels is typically large enough to produce sparse expected contingency tables. A direct consequence of this is that $\operatorname*{L_{\text{max}}}$ becomes very large, leading to potentially redundant calculations and therefore an increased computational cost. This is observed in some cases in our simulations, where the depth is much lower than the computed value of $\operatorname*{L_{\text{max}}}$, indicating that sequences of greater length have limited to no contribution to the obtained scores of nominal outlyingness. An amendment in the way $\operatorname*{L_{\text{max}}}$ is determined can alleviate this issue and reduce the computational complexity, without changing the way the results are interpreted and the conclusions derived. Additionally, one may wish to extend the simulation study by enforcing a dependence structure among variables which are likely to be associated in a specific manner. This way, we can get rid of the independence assumption (which is typically not valid) and potentially obtain more interesting results.
@@
We conclude by giving some recommendations on how our framework may be used for extending the research that is currently being conducted in certain fields. The proposed framework can be used for assigning weights to observations in a data set prior to running a classification algorithm. As an example, one may perhaps want to downweigh subjects with a large score of nominal outlyingness, as these can be seen as potential outliers which may distort the classification output. This can be done by modifying the relevant loss function to account for observation weights. Similarly, one might wish to implement such an approach for clustering categorical data or deriving a version of the Trimmed K-Means [17] algorithm for nominal data. Finally, the proposed framework can contribute in the development of more reliable predictions in linear regression with nominal predictors by diminishing the impact of potential outliers, yielding robust regression algorithms for nominal data. These recommendations can pave the way for future research opportunities and culminate in the development of novel methodologies which in turn provide insight on research questions of interest.
@@
The R code for the simulations conducted is available on GitHub at https://github.com/EfthymiosCosta/SONO_Sims.